# Autoregressive Stochastic Clock Jitter Compensation in Analog-to-Digital Converters

### v2

@@ I Introduction
Analog-to-Digital Converters are well-known to introduce a certain number of distortions in the sampled signals; this is usually due to hardware imperfections, and they manifest as DC offsets biases, nonlinear distortions due to clipping and irregularities in the sampling instants (jitter) [36].
The ADC clock normally consists of a voltage controlled oscillator (VCO), which is phase locked to a reference clock with a certain loop bandwidth. The phase locked loop (PLL) noise within the loop bandwidth normally has a scaled relationship to the reference phase noise whereas the noise outside the loop bandwidth is related to the VCO and the clock distribution noise. This phase noise, when converted to the time domain, manifests as clock jitter, meaning that the cumulative phase fluctuations from both the PLL-regulated and the free-running portions of the VCO directly contribute to time jitter, which ultimately affects the timing accuracy and performance of the ADC system.
In a noisefree situation, perfect signal recovery when the sampling is non-uniform but deterministic can be done via the classical Paley-Wiener-Levinson theory [21]. Prompted by the accelerating technical development of ADCs, research tackling the problem of random timing offsets has been performed already since the early $60$s [6][4]; in this context the target optimality is not pointwise, but rather some mean squared error; see [37][27][12].
Some research has been devoted to jitter identification and compensation in modern ADCs, for instance via pilot signals injection [33][29][35]. When pilot sinusoids are utilized, the jitter manifests itself as phase noise disturbance and therefore phase noise-dedicated techniques can be employed to track and compensate it [33][29][35]. Despite being often and widely modeled as uncorrelated additive Gaussian uncertainty [37][34][3][16], it is well known that the clock jitter process does not have a white spectrum nor can be usually considered uncorrelated [7]; moreover partially contradicting models for the jitter and its power spectral density (PSD) are sometimes found in literature - see e.g. [35] as opposed to [8][39], and how they differently model it near 00 Hz. There is however a widespread agreement on its $1/f^{2}$ spectral decay in mid-band decades, that shows up in all the mathematical models surveyed. The latter property, together with correlated behavior to describe “slow” and “fast” jitter [7] can be compactly captured by means of autoregressive (AR) processes, or more generally by autoregressive moving-average (ARMA) or autoregressive fractionally integrated moving-average (ARFIMA) models, whose spectral densities could be even more flexibly prescribed so that even the narrow near-DC close-in $1/f^{3}$ region [39] could be reproduced by adding additional ARMA poles. In the present we thus follow the autoregressive model proposed in [7].
@@
    rigorous justifications for the approximations arising in Taylor expansions and outline of the regime in which these approximations are valid (Section II-D);
@@ I-A Notation and symbols
ADCs act on continuous-time signals by discretizing and digitalizing them, so that the output of an ADC is a discrete sequence of values $x(t_{n})$ when the input is the analog $x(t)$. If $x(t)$ is a stochastic process, we can model the ADC output as a sequence of random variables; we will indicate sequences (finite or infinite) with bold lowercase letters $\mathbf{x}\coloneqq(x_{1},x_{2},\dots)$; if $\mathbf{x}$ and $\mathbf{y}$ are two such sequences, $\mathbf{x}\odot\mathbf{y}$ indicates their Hadamard componentwise product. Bold uppercase letters $\mathbf{A},\mathbf{B},\dots$ will be reserved to matrices; the symbol $\approx$” stands for “approximately equal to” (with some negligible error) while the symbol $\ll$” means “much less than”, more precisely $x\ll y$ if and only if $x/y<K$ with $K\leq 1/10^{2}$. With $\widehat{\cdot}$ we will indicate estimations of sought quantities, and with $\widetilde{\cdot}$ noisy measurements of quantities.
@@
Oftentimes throughout the text we will encounter continuous-time processes $y$ (Wiener, Ornstein-Uhlenbeck, white noise) whose derivative might not exist anywhere in the classical sense. However their discrete $D\mathbf{y}$ (bandpassed, and possibly aliased) counterparts will exist; to keep the notation as light as possible we will write, with a slight abuse of notation, $y^{\prime}_{n}$ in lieu of $(D\mathbf{y})_{n}$.
II System model
II-A Signal model
In the present work, we consider baseband analog signals that are modeled as continuous-time bandlimited real-valued stationary Gaussian processes $x(\omega,t)$ so that, for each time instant $t\in T$, the random variable $(\omega\mapsto x(\omega,t))\sim\mathcal{N}(0,1)$; $x$ has flat power spectral density $\mathcal{S}_{x}(f)=1/(2W)\chi_{\{|f|\leq W\}}(f)$, where $\chi(\cdot)$ denotes the indicator function. This is a standard model used to describe signals transmitted in telecommunication applications, cfr. [25]. We recall that sample paths of these processes are analytically well-behaving, i.e. the maps $t\mapsto x(\omega,t)$ are entire (and thus differentiable with continuity) for almost every $\omega\in\Omega$, as direct consequence of e.g. Theorem 11 in [5]. Therefore the objects $x^{(n)}$ are well-defined for all $n$.
II-B Stochastic jitter description and problem statement
@@
$\mathcal{S}_{\xi}(e^{i\omega})=\frac{\sigma_{\epsilon}^{2}}{1+\varphi^{2}-2\varphi\cos(\omega)},\ \omega\in[-\pi,\pi]$ (2)
from which it is pretty straightforward to observe that much of the process power is concentrated near $\omega=0$ (or 00 Hz) when $\varphi\approx 1$:
$\mathcal{S}_{\xi}(e^{i\omega})\approx\frac{\sigma_{\epsilon}^{2}}{(1-\varphi)^{2}+\varphi\omega^{2}};$ (3)
in fact, the latter is the PSD of the Ornstein-Uhlenbeck process, that might be regarded as the continuous counterpart of an AR(1). The AR(1) jitter description formulation offers a faithful (especially in describing the quadratic spectral decay (2), cfr. [9]), convenient and mathematically tractable representation of the disturbance process, which appears to have been overlooked in the literature.
@@
As we show later in Proposition II.1, the higher order terms in the Taylor expansion can be considered negligible. From here on, we will also set
@@
as our model under investigation, with $w_{n}\sim\mathcal{N}(0,\sigma_{w}^{2})$ a sequence of independent Gaussian random variables, which are also independent of the process $\{\xi_{n}\}_{n\in\mathbb{N}}$.
@@
II-C How much jitter is “too much” jitter?
@@
$\begin{split}\text{SINADR}_{\text{dB}}&=10\log_{10}\left[\frac{\text{Var}(x_{n})}{\text{Var}(w_{n})+\text{Var}(\xi_{n}x^{\prime}_{n})}\right]\\ &=10\log_{10}\left[\frac{\text{Var}(x_{n})}{\text{Var}(w_{n})+\text{Var}(\xi_{n})\text{Var}(x^{\prime}_{n})}\right]\\ &=10\log_{10}\left[\frac{3(1-\varphi^{2})}{3(1-\varphi^{2})\sigma_{w}^{2}+4\pi^{2}W^{2}\sigma_{\epsilon}^{2}}\right],\end{split}$ (6)
where we used $\text{Var}(\xi_{n})=\sigma_{\epsilon}^{2}/(1-\varphi^{2})$, $\mathbb{E}[\xi_{n}]=\mathbb{E}[x^{\prime}_{n}]=0$ and the fact that $\xi_{n}$ and $x_{n}$ are independent. By means of inverting eq. (6) it is easy to see how much jitter could be tolerated (for a desired SINADR level). At the same time for the first-order Taylor expansion (4) to hold, the jitter $\xi_{n}$ cannot be too large; therefore one reasonable assumption is to make the “small jitter” hypothesis (in line with the literature surveyed so far), i.e. to request
@@
we will refer to the left hand side of the latter as jitter percentage.
II-D Approximation errors analysis
There are two main approximations that will be performed in our analysis, whose errors need to be rigorously analyzed. In the first-order expansion in (4) under (II-C) the neglected higher-order terms are, in some sense, negligible, as the following Proposition clarifies:
@@ II-D Approximation errors analysis
Let $\mathbf{x}=\{x_{n}\}_{n\geq 1}$ be a discrete zero-mean, wide-sense Gaussian stationary bandlimited process and $\boldsymbol{\xi}=\{\xi_{n}\}_{n\geq 1}$ a discrete autoregressive process of order $1$ with variance $\sigma_{\xi_{n}}^{2}\equiv\sigma_{\xi}^{2}$, independent on $\mathbf{x}$. Assume that (II-C) holds; then we have
@@
Moreover
$R(x_{n},\xi_{n})\coloneqq\frac{\mathbb{E}[(x(\cdot,\ nT_{s}+\xi_{n})-x_{n}-x_{n}^{\prime}\xi_{n})^{2}]}{\mathbb{E}[(x_{n}+x_{n}^{\prime}\xi_{n})^{2}]}\ll 1.$ (7)
@@
where we set $\mathbf{u}=\boldsymbol{\xi}\odot D\mathbf{x}$. The term $w^{\prime}_{n}$ is violet noise (bandlimited to $1/2T_{s}$), whose variance equals $\pi^{2}\sigma_{w}^{2}/3T_{s}^{2}$. For a fixed Noise-to-Distortion Ratio (NDR) $r$ (in dB), the white noise variance satisfies $\sigma_{w}^{2}=10^{r/10}\text{Var}(\xi x^{\prime})$. Therefore
@@
and thus $\sigma_{w^{\prime}}^{2}\ll\sigma_{x^{\prime}}^{2}$ provided that (II-C) holds and the white noise is not too strong. The term $D(\boldsymbol{\xi}\odot\mathbf{x}^{\prime})_{n}$ has a more convoluted structure, but its impact can be precisely gauged, as shown in Proposition II.2, whose proof can be found in the appendix. Figure 1 shows again that $\sigma_{x^{\prime}}^{2}$ is fully dominating.
@@ III Pilot samples-based jitter compensation
In this section are found the main contributions of this paper. We present two jitter tracking techniques aimed at mitigating the negative impact of this type of disturbance. Both algorithms will assume the knowledge of pilot samples at known time index positions, meaning that the values of $x_{n}$ for $n\in I\subseteq\mathbb{N}$ (for example so that the pilots are uniformly spaced) will be assumed to be known a priori. This can be achieved for example by time-multiplexing known values to the sampled unknown signal.
A noisy estimation of the jitter process sample values can be obtained by reversing equations (4) and (5) at pilot positions:
@@
As opposed to pilot tones-based techniques [35], we shall note that no knowledge of the signal’s derivative is assumed by our techniques, i.e. $x^{\prime}_{n}$ is not a priori known at pilot positions (nor anywhere else). Moreover, the knowledge of the signal’s samples $x_{n}$ at pilot positions is not enough to compute its derivative at these positions. In our algorithms we will use $y^{\prime}_{n}$ instead, whose utilization is justified by the analysis carried out in Section II-D.
Figure 2: Jitter tracker comparison example, Algorithm 1 vs Algorithm 2.
@@ III-A Robust blockwise polynomial interpolation
The hidden process is treated here as blockwise deterministic (yet unknown), uncorrelated and slowly varying; under the model (4), noisy estimations of the jitter samples values at pilot positions can be obtained from (11) as $\widetilde{\xi}_{n}=\xi_{n}+w_{n}/x_{n}^{\prime}$. The additive noise is conditionally Gaussian but heteroscedastic because its variance is time-dependent; moreover the (unknown) values $x^{\prime}_{n}$ are to be replaced with their estimations $y^{\prime}_{n}$. We set $v_{n}=w_{n}/y^{\prime}_{n}$. The idea is to perform an optimal short-time ($=$ blockwise) polynomial interpolation of the jitter process at query points $\widetilde{\xi}_{n}$, $n\in I$ for a certain $I$.
@@
$\widehat{\xi}_{n}=\begin{cases}p_{1}(nT_{s}),&1\leq n\leq i_{C}^{(1)}\\ p_{l}((n-i_{1}^{(l)}+1)T_{s}),&i_{1}^{(l)}<n\leq i_{C}^{(l)},\ l=2,\dots,L;\end{cases}$ (16)
notice that the latter minimizes the conditional mean-square error $\mathbb{E}[(\widehat{\xi}_{n}-\xi_{n})^{2}\,|\,\widetilde{\mathbf{y}}]$ at least at $n\in I$.

### v3

@@ I Introduction
Analog-to-digital converters introduce various distortions into sampled signals. These arise from hardware imperfections and appear as DC offset biases, nonlinear distortion from clipping, and irregularities in sampling instants (jitter) [43].
The ADC clock typically uses a voltage controlled oscillator locked to a reference clock by a phase locked loop with a defined loop bandwidth. Inside that bandwidth, loop noise relates in scale to the reference phase noise; outside it, noise originates from the oscillator itself and the clock distribution network. When converted to time, this phase noise appears as clock jitter. Cumulative phase fluctuations from both the loop regulated portion and the free running oscillator directly add to time jitter, degrading the timing accuracy and overall performance of the converter.
In the absence of noise, perfect recovery of a deterministic but non uniform sampling pattern can be achieved using classical Paley-Wiener-Levinson theory [25]. Driven by advances in converter technology, work on random timing offsets began in the early 1960s [8][4]. In that setting optimality is measured not at each sample point but in terms of mean squared error [44][32][15].
Modern work on identifying and compensating jitter in converters has included techniques based on injecting pilot tones [40][34][42]. When pilot sinusoids are present, jitter appears as phase noise disturbance, allowing dedicated phase noise tracking and compensation methods to be applied [40][34][42]. Although jitter is often modeled as uncorrelated additive Gaussian uncertainty [44][41][3][20], the clock jitter process neither has a white spectrum nor is generally uncorrelated [9]. Conflicting models of its power spectral density appear in the literature - for example [42] versus [11][46] -particularly in their treatment near zero Hertz. However, all surveyed models agree on a $1/f^{2}$ spectral decay in mid-band decades. That property, along with correlated behavior capturing slow and fast jitter [9], can be modeled by autoregressive (AR) processes or more generally by autoregressive moving-average (ARMA) processes. Experiments reported in [46] show a narrow region near DC exhibiting a $1/f^{3}$ decay. Capturing that requires more flexible models such as autoregressive fractionally integrated moving average (ARFIMA) processes, which introduce additional complications from potential non-stationarity. Standard Kalman filtering must also be adapted to handle state-space models with fractional operators [39].
@@
    rigorous justifications for the approximations arising in Taylor expansions and outline of the regime in which these approximations are valid (Section II-C);
@@ I-A Notation and symbols
ADCs act on continuous-time signals by discretizing and digitalizing them, so that the output of an ADC is a discrete sequence of values $x(t_{n})$ when the input is the analog $x(t)$. If $x(t)$ is a stochastic process, we can model the ADC output as a sequence of random variables; we will indicate sequences (finite or infinite) with bold lowercase letters $\mathbf{x}\coloneqq(x_{1},x_{2},\dots)$; if $\mathbf{x}$ and $\mathbf{y}$ are two such sequences, $\mathbf{x}\odot\mathbf{y}$ indicates their Hadamard componentwise product. Bold uppercase letters $\mathbf{A},\mathbf{B},\dots$ will be reserved to matrices; the symbol $\approx$” stands for “approximately equal to” (with some negligible error) while the symbol $\ll$” means “much less than”. With $\widehat{\cdot}$ we will indicate estimations of sought quantities, and with $\widetilde{\cdot}$ noisy measurements of quantities.
@@
Oftentimes throughout the text we will encounter continuous-time processes $y$ (Wiener, Ornstein-Uhlenbeck, white noise) whose derivative might not exist anywhere in the standard pointwise sense. However, upon bandlimiting and sampling them, their discrete $D\mathbf{y}$ counterparts will exist; to keep the notation as light as possible we will write, with a slight abuse of notation, $y^{\prime}_{n}$ in lieu of $(D\mathbf{y})_{n}$.
II Signal and system model
We consider baseband analog signals that are modeled as continuous-time bandlimited complex-valued stationary Gaussian processes $x(\omega,t)$ so that, for each time instant $t\in T$, the random variable $(\omega\mapsto x(\omega,t))\sim\mathcal{N}(0,1)$; $x$ has flat power spectral density $\mathcal{S}_{x}(f)=1/(2W)\chi_{\{|f|\leq W\}}(f)$, where $\chi(\cdot)$ denotes the indicator function. This is a standard model used to describe signals transmitted in telecommunication applications, cfr. [30]. We recall that sample paths of these processes are analytically well-behaving, i.e. the maps $t\mapsto x(\omega,t)$ are entire (and thus differentiable with continuity) for almost every $\omega\in\Omega$, as direct consequence of e.g. Theorem 11 in [5]. Therefore the “n-th derivatives objects” $x^{(n)}$ are well-defined for all $n\in\mathbb{N}$.
II-A Stochastic jitter description and problem statement
@@
$\mathcal{S}_{\xi}(e^{i\omega})=\frac{\sigma_{\epsilon}^{2}}{1+\varphi^{2}-2\varphi\cos(\omega)},\ \omega\in[-\pi,\pi],$ (2)
from which it is pretty straightforward to observe that much of the process power is concentrated near $\omega=0$ (or 00 Hz): indeed
$\begin{split}\mathcal{S}_{\xi}(e^{i\omega})&=\frac{\sigma_{\epsilon}^{2}}{(1-\varphi)^{2}+2\varphi(1-\cos(\omega))}\\ &=\frac{\sigma_{\epsilon}^{2}}{(1-\varphi)^{2}+4\varphi\sin^{2}(\omega/2)}\\ &\leq\frac{\sigma_{\epsilon}^{2}}{(1-\varphi)^{2}+4\varphi\omega^{2}/\pi^{2}}\end{split}$ (3)
where we used $\sin(\omega)\geq 2\omega/\pi$ for $\omega\in[0,\pi/2]$, from which it is again evident the desired quadratic decay in $\omega$. In fact, the last formula in (3) is the PSD of the Ornstein-Uhlenbeck process, that might be regarded as the continuous counterpart of an AR(1). The AR(1) jitter description formulation offers a faithful, convenient and mathematically tractable representation of the disturbance process, which appears to have been overlooked in the literature.
@@
As we show later in Proposition II.1, the higher order terms in the Taylor expansion can be considered in some sense negligible.
From here on, we will also set
@@
as our model under investigation, with $w_{n}\sim\mathcal{N}(0,\sigma_{w}^{2})$ a sequence of independent Gaussian random variables modeling additional white noise in measurements, which are also independent of the process $\{\xi_{n}\}_{n\in\mathbb{N}}$.
@@
II-B How much jitter is “too much” jitter?
@@
$\begin{split}\text{SINADR}_{\text{dB}}&=10\log_{10}\left[\frac{\text{var}(x_{n})}{\text{var}(w_{n})+\text{var}(\xi_{n}x^{\prime}_{n})}\right]\\ &=10\log_{10}\left[\frac{\text{var}(x_{n})}{\text{var}(w_{n})+\text{var}(\xi_{n})\text{var}(x^{\prime}_{n})}\right]\\ &=10\log_{10}\left[\frac{3(1-\varphi^{2})\sigma_{x}^{2}}{3(1-\varphi^{2})\sigma_{w}^{2}+4\pi^{2}W^{2}\sigma_{\epsilon}^{2}\sigma_{x}^{2}}\right],\end{split}$ (6)
where we used $\sigma_{\xi}^{2}=\sigma_{\epsilon}^{2}/(1-\varphi^{2})$, $\mathbb{E}[\xi_{n}]=\mathbb{E}[x^{\prime}_{n}]=0$ and the fact that $\xi_{n}$ and $x_{n}$ are independent. By means of inverting eq. (6) it is easy to see how much jitter could be tolerated (for a desired SINADR level). At the same time for the first-order Taylor expansion (4) to hold, the jitter $\xi_{n}$ cannot be too large; therefore one reasonable assumption is to make the “small jitter” hypothesis (SJH), which is in line with the literature surveyed so far, i.e. to request
@@
we will refer to the left hand side of the latter as jitter percentage. Notice that when $\sigma_{w}=0$, (6) becomes equal to
$-20\log_{10}\left[2\pi W\sigma_{\xi}/\sqrt{3}\right]\approx-20\log_{10}\left[2\pi W\sigma_{\xi}\right]+4.77\text{ dB},$
which is the standard jitter SNR formula found in classical literature (see e.g. [36]), applied to bandlimited signals.
The interaction between the two additive distortions in (5) will also play a role in our analysis; therefore we introduce a convenient notion of Noise-to-Distortion Ratio (NDR), an obvious variant of (6):
$\text{NDR}_{\text{dB}}=10\log_{10}\left[\frac{\sigma_{w}^{2}(1-\varphi^{2})}{4\pi^{2}W^{2}\sigma_{x}^{2}\sigma_{\epsilon}^{2}}\right].$ (6)
II-C Approximation errors analysis
There are two main approximations that will be performed in our analysis, whose errors need to be rigorously analyzed. In the first-order expansion in (4) under (II-B) the neglected higher-order terms are, in some sense, negligible, as the following Proposition clarifies:
@@ II-C Approximation errors analysis
Let $\mathbf{x}=\{x_{n}\}_{n\geq 1}$ be a discrete zero-mean, wide-sense Gaussian stationary bandlimited process and $\boldsymbol{\xi}=\{\xi_{n}\}_{n\geq 1}$ a discrete autoregressive process of order $1$ with variance $\sigma_{\xi_{n}}^{2}\equiv\sigma_{\xi}^{2}$, independent on $\mathbf{x}$. Assume that (II-B) holds; then we have
@@
where we set $\mathbf{u}=\boldsymbol{\xi}\odot D\mathbf{x}$. The term $w^{\prime}_{n}$ is violet noise (bandlimited to $1/2T_{s}$), whose variance equals $\pi^{2}\sigma_{w}^{2}/3T_{s}^{2}$. For a fixed NDR level $r$, the white noise variance satisfies $\sigma_{w}^{2}=10^{r/10}\text{var}(\xi x^{\prime})$. Therefore
@@
and thus $\sigma_{w^{\prime}}^{2}\ll\sigma_{x^{\prime}}^{2}$ provided that (II-B) holds and the white noise is not too strong. The term $D(\boldsymbol{\xi}\odot\mathbf{x}^{\prime})_{n}$ has a more convoluted structure, but its impact can be explicitly assessed, as shown in Proposition II.2, whose proof can be found in the appendix. Figure 1 shows again that $\sigma_{x^{\prime}}^{2}$ is fully dominating.
@@
Remark II.1.
We point out that a stochastic version of the Peano remainder theorem for Taylor expansions hold, cfr. Lemma 3 in [45]. Further results comparing (conditional) variances of the objects $x(\omega,nT_{s}+\xi_{n})$ and $x(\omega,nT_{s})+x^{\prime}(\omega,nT_{s})\xi_{n}$ would require some care; first of all the measurability/randomness of the maps $\omega\mapsto x(\omega,nT_{s}+\xi_{n}(\omega))$ or $(\omega_{1},\omega_{2})\mapsto x(\omega_{1},nT_{s}+\xi_{n}(\omega_{2}))$ must be carefully assessed, depending on whether the jitter process is seen as living in the same probability space. Measurability of $x(\omega,t)$ on the whole product space $\Omega\times\mathbb{R}$ typically require left- or right-continuity in $t$ to hold for all $\omega\in\Omega$ (see Lemma 6.4.6 in [7]), which might not be the case for our signal model. In that case one shall replace $x$ with an indistinguishable copy for which the continuity holds for every $\omega\in\Omega$.
@@ III Pilot samples-based jitter compensation
In this section are found the main contributions of this paper. We present two jitter tracking techniques aimed at mitigating the negative impact of this type of disturbance. Both algorithms will assume the knowledge of pilot samples at known time index positions, meaning that the true sample values at $n\in I\subsetneq\mathbb{N}$ will be assumed to be known a priori. This may achieved be programming the analog front end so that at specific time instants the ADC is fed with a known analog stimulus; the outputted stream of samples will be jittered, including the samples at pilot positions. A noisy estimation of the jitter process sample values can be obtained by reversing equations (4) and (5) at pilot positions:
@@
Once the jitter process is estimated using one of the methods detailed in this section, the jitter distortion in (5) must be subtracted, the pilots removed and the signal interpolated using e.g. the Gerchberg-Papoulis algorithm [27][18] to “fill the holes” left after the removal of pilot samples. The workflow is summarized by the scheme in Figure 2; the focus of this paper is within the green box.
Introduce known analog stimulus. Use pilot samples to estimate $\xi_{n}\,\forall\,n$ and compensate the distorted signal. Remove pilot samples and interpolate (w e.g. Gerchberg-Papoulis).
Figure 2: Schematic three‐step process.
As opposed to pilot tones-based techniques [42], we shall note that no knowledge of the signal’s derivative is assumed by our techniques, i.e. $x^{\prime}_{n}$ is not assumed to be a priori known at pilot positions (nor anywhere else). Moreover, the knowledge of the signal’s samples $x_{n}$ at pilot positions is not enough to compute its derivative at these positions. In our algorithms we will use $y^{\prime}_{n}$ instead, whose utilization is justified by the analysis carried out in Section II-C. An example of tracking is showcased in Figure 3; despite the large amount of missing data, both methods are able to reconstruct the aggregated behaviour of the underlying autoregressive process.
Figure 3: Jitter trackers comparison example, Algorithm 1 (left) vs Algorithm 2 (right).
@@ III-A Robust blockwise polynomial interpolation
The hidden true process is treated here as blockwise deterministic (yet unknown), uncorrelated and slowly varying; under the model (4), noisy estimations of the jitter samples values at pilot positions can be obtained from (11) as $\widetilde{\xi}_{n}=\xi_{n}+w_{n}/x_{n}^{\prime}$. The additive noise is conditionally Gaussian but heteroscedastic because its variance is time-dependent; moreover the (unknown) values $x^{\prime}_{n}$ are to be replaced with their estimations $y^{\prime}_{n}$. We set $v_{n}=w_{n}/y^{\prime}_{n}$. The idea is to perform an optimal short-time ($=$ blockwise) polynomial interpolation of the jitter process at query points $\widetilde{\xi}_{n}$, $n\in I$ for a certain $I$.
@@
$\widehat{\xi}_{n}=\begin{cases}p_{1}(nT_{s}),&1\leq n\leq i_{C}^{(1)}\\ p_{l}((n-i_{1}^{(l)}+1)T_{s}),&i_{1}^{(l)}<n\leq i_{C}^{(l)},\ l=2,\dots,L.\end{cases}$ (16)